Subject: APO 3.5m Users committee minutes, 2/2/09

From: Suzanne Hawley

Submitted: Fri, 6 Feb 2009 10:51:33 -0800 (PST)

Message number: 1125 (previous: 1124, next: 1126 up: Index)

               APO 3.5-m Users Committee Phonecon, 2/2/09

Attending:  Suzanne Hawley, Scott Anderson, Mark Klaene, Bill 
Ketzeback, John Bally, Remy Indebetouw, Bruce Gillespie, Michael 
Strauss, Al Harper, Jon Holtzman, Cynthia Froning, and Jon Fulbright

Absent: none


User feedback, comments from institutional representatives:

Princeton (Michael Strauss) - Michael said there were no concerns. 
Princeton astronomers are currently at APO for an observing run.
Suzanne appreciated their help in accomodating a late schedule change.

Colorado (John Bally) - John reported that John Stocke recently used 
the new NIC-FPS filter focus offsets, and they seemed to work well. 
Suzanne added we should now be using these offsets routinely, and to 
make sure that they are well-documented in the users manual.

Johns Hopkins (Jon Fulbright) - Jon had nothing to report.

Washington (Scott Anderson) - Scott had nothing to report.

Chicago (Al Harper) - Al had nothing to report, other than to remark 
that the UC astronomers are interested in the outcome of the tertiary 
pointing problems and the focus offsets for the NIC-FPS narrow-band 

New Mexico State (Jon Holtzman) - Jon had nothing to report.

Virginia (Remy Indebetouw) - Remy had nothing to report.


Discussion of telescope/instruments report - Klaene:

Mark pointed to the good weather (albeit with variable seeing) that 
we've enjoyed during the past month.  The principal telescope issue 
has been pointing problems, of around 3 arcminutes, that are suspected 
to be related to the tertiary mirror mount.  Significant engineering 
work has gone into investigating the problem so far, and much of next 
week will be spent trying to track down and fix the problem.  In the 
meantime, operational workarounds have been effective in minimizing 
down time.  The instruments are all working relatively well.  We're 
still progressing on getting CU/UVa to replace the NIC-FPS fan-out 
board (John Bally will check status).  Agile is ready to move to the 
new side port on the telescope, but this work has been held up by the 
problem with the tertiary mirror mount.


3.5-m Telescope, Instruments, and CIF Projects Highlights, 1/8/09 
through 1/28/09

                                Mark Klaene

0) Overview

Telescope problems were generally associated with pointing and the 
tertiary.  Instruments were operational this period.  Generally clear 
skies prevailed with moderate temperatures.

1) Telescope

NA2 axis controller continues to perform well.

Tertiary rotation: The tertiary is now rotating well, but when it 
arrives at a port, there have been issues with the mirror tilts being 
applied.  Multiple problems with the B actuator were uncovered. One of 
the problems was mechanical that prompted a replacement of the 
actuator itself.  Another intermittent problem is also occurring, 
causing significant pointing errors. This is still under investigation 
since we can not seem to get it to repeat while troubleshooting during 
the day.  We will be carrying out additional work on the tertiary 
during the February engineering run.

2) Instruments

DIS, SPIcam, and Echelle are fully operational.

NIC-FPS is operational with no problems.  We are still waiting for UVa 
to design and build up a new NIC-FPS fan-out board that will hopefully 
address the residual readout noise.  Currently UVa and CU are 
reviewing the design.

TripleSpec mounting is still occasionally giving us some problems, but 
recent modifications and more training seem to have improved the 
situation.  The current state is more problematic than we would like, 
but the situation is tolerable at the moment.

Agile is still available at NA2. TUI timing tests are in work to 
verify precision timing.  Rotator testing has been delayed in order to 
focus on the tertiary issues.

APOLLO remains operational.  The high-precision gravimeter is now 
slated for arrival in February, and preparations are in progress for 
its installation at the 3.5-m.

3) CIF projects

Preparation continues for an early March direct-drive CDR.


New/upgraded instrument plans - Hawley

This topic is about the planning process to upgrade our direct visible- 
imaging capability, the echelle, and a newly discussed proposal to 
upgrade the Goddard Fabry-Perot (GFP) after recent discussions at the 
site with Bruce Woodgate.  On the latter, Bruce Woodgate is planning 
to upgrade the GFP detector to improve sensitivity, and add an 
integral-field unit to its optics, which will afford IFU capability 
for spectroscopy over a 7 or 14 arcsecond field.  We want to hear if 
the user community would find the GFP more generally useful for 
science if these improvements are made.  Bruce Woodgate and Carol 
Grady have also offered to run a GFP training class for potential users.
Suzanne has asked Bruce Woodgate to write a white 
paper on the GFP upgrades that we can send around to the general user 
community--we'll discuss this further at the next Users Committee 
meeting.  In the meantime, Suzanne asked the Users Committee members 
to think about the GFP and upgrades and talk to their users about it.

Suzanne added that Eric Burgh (CU) is leading the planning to upgrade 
the echelle; a memo was sent around last week to the users to provoke 

Cynthia Froning (CU, with John Bally) is looking at upgrading or 
replacing SPIcam.  The trades between field-of-view and specialized 
capabilities are being explored.  She plans to send a general call to 
the users asking for input, and in parallel will form a small working 
group to define priorities and help lead the design and 
implementation.  Suzanne said that 1 or 2 interested people from each 
institution should join this working group, and asked that the Users 
Committee member each nominate people from their institutions and send 
their names to Cynthia (cynthia.froning at, copied to 

Note: there are two other white papers for new instruments presently
in circulation, both from Jian Ge at U. Florida. One is for a
high resolution NIR spectrograph, the other for a NIR Doppler
interferometer for precision radial velocities.  Both would
use fiber feeds from the echelle port.  Feedback on these ideas
is also welcome.

Users committee reps are asked to discuss all of these instrument
options with their communities.


IYA at APO - Klaene

Make reported that for the International Year of Astronomy, APO will 
be featured in a 20-minute segment during a 100-hour webcast on April 
3.  On April 10, a half night of engineering time will be used to host 
a public open house at the 3.5-m.  Bill asked if Adler had any 
interest in participating, and in restarting their Friday-night Adler 
public program about APO.  Al said he would ask, and report next month.


Decadal Survey input - Hawley

There are two Decadal Survey groups within infrastructure studies
that have overlap with  topics of interest for the 3.5-m (partnerships, 
funding, private and moderate-aperture telescopes).  There may 
be opportunities to contribute white papers on these topics to the 
study groups, in the context of our situation and plans.  Suzanne 
asked that all users give her input on what issues they feel need 



[open from previous months]:

ACTION:  Focus offsets for the various NIC-FPS filters need to be 
determined, documented for the users, and automated corrections 
implemented in the operations software.

    STATUS:  Closed.  We have posted the offset numbers for NIC-FPS 
filter focus offsets on the website.  For now, users should
check with the Observing Specialists to insure that the offsets are
being applied.  Another set of offset measurements will be made as a lower 
priority in monthly engineering time, and if users want to contribute 
some of their time for these measurements, they should talk to Bill 
Ketzeback or Russet McMillan.  The issue of applying automatic offsets via 
instrument and telescope control software will be added to the list 
of engineering projects.  John Bally will consult with Stephane Beland
about the role of the NIC-FPS ICC.

[new actions from this meeting]:



Next meeting: The next Users Committee phonecon will be on Monday, 2 
March, at 8:30 AM Pacific Time.  The agenda and other materials will 
be sent to the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1125 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to